# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v1

@@ Appendix E Limitations
Thermalization. This work depends on the thermalization of the final renders in the FAKE dataset. The dataset contains very difficult lighting conditions and scene compositions that make it powerful to train landmarkers but also made the thermalization particularly challenging and could only be solved with advanced regularization approaches. Despite a good perceptual result, some T-FAKE images contain artifacts, e.g., black patches on the upper body. Nevertheless, these artifacts remain limited.
Dense Landmarks. In this work, we relied on the ground truth 70-point landmarks of the FAKE dataset [52] and dense Mediapipe [35] annotations. Training with the original 320-point and 702-point landmarks could further boost accuracy and lead to an even denser landmarker. However, these landmarks are not publicly available. Furthermore, we only evaluate a mobilenet backbone. Performance for difficult poses and face variability could be achieved with denser models [53].
CHARLOTTE. The CHARLOTTE [3] dataset is among the largest datasets with thermal recordings of faces that contain different levels of image quality as well as has a high variability in poses such as side profile pictures and tilting which makes the dataset ideal as a benchmark. However, the annotation uses a non-standard convention where side profile images have a different number of landmarks than frontal faces. Furthermore, landmarks of low-resolution images are quantized and there are examples of shifted ground truth annotations (see Figure 9).
@@
Label Adaptation. We retrain the label adaptation networks for each tested landmarker on its original predictions. Hence, for a given method, a high failure rate on CHARLOTTE means that fewer training images are available for that method. Also, very poor predictions on some of the images such as profile images without failure produce a low quality of adapted landmarks (see Figure 10, TL54 dlib, column 3). It is important to note that the label adaptation does not use visual information from the benchmark dataset to translate landmarks. However, the same individuals were present in both the test and the training dataset splits which might allow the network to learn the facial statistics of individuals. This might give an advantage to landmarkers with only a few landmarks, e.g., TFW [32] (see Figure 10, TFW). The high 68-point landmark performance of TFW on the CHARLOTTE thermal data does not necessarily mean that this generalizes and the advantage of increasing the number of predicted landmarks has been demonstrated for RGB images [53].

### v2

@@ Appendix E Limitations
Thermalization. This work depends on the thermalization of the final renders in the FAKE dataset. The dataset contains very difficult lighting conditions and scene compositions that make it powerful to train landmarkers but also made the thermalization particularly challenging and could only be solved with advanced domain-adaptive semi-supervised regularization approaches. Despite a good perceptual result of the faces, some T-FAKE images can contain minor artifacts on clothing and on the background (e.g. see Figure 12, bottom second from right). Nevertheless, these artifacts remain limited. Moreover, background artifacts can easily be removed by choosing a suitable background based on the ground truth segmentation as implemented during our landmarker training, see Fig. 18. We merely used an MSE loss for our supervised training. Including an adversarial [26, 64] or a perceptual [27, 53] loss into our model should lead to additional perceptual improvement.
Dense Landmarks. In this work, we relied on the ground truth 70-point landmarks of the FAKE dataset [61] and dense Mediapipe [40] annotations. Training with the original 320-point and 702-point landmarks could further boost accuracy and lead to an even denser landmarker. However, these landmarks are not publicly available. Furthermore, we only evaluate a mobilenet backbone for landmark detection without face tracking. Better performance for difficult poses and face variability could be achieved with denser models [62] and spatial normalization of face positions during training.
CHARLOTTE. The CHARLOTTE [3] dataset is among the largest datasets with thermal recordings of faces that contain different levels of image quality as well as has a high variability in poses such as side profile pictures and tilting which makes the dataset ideal as a benchmark. However, the 2D annotation uses a convention where side profile images have a different number of landmarks than frontal faces. Furthermore, landmarks of low-resolution images are quantized and there are examples of shifted ground truth annotations (see Figure 14).
@@
Label Adaptation. We retrain the label adaptation networks for each tested landmarker on its original predictions. Hence, for a given method, a high failure rate on CHARLOTTE means that fewer training images are available for that method. Also, poor predictions on some of the images such as profile images without failure produce a low quality of adapted landmarks (see Figure 15, TL54 dlib, column 3). It is important to note that the label adaptation does not use visual information from the benchmark dataset to translate landmarks. However, the same individuals were present in both the test and the training datasets which might allow networks to learn facial statistics of individuals. This might give an advantage to landmarkers with only a few landmarks, e.g., TFW [37] (see Figure 15, TFW). The good 68-point landmark performance of TFW on the CHARLOTTE dataset does not necessarily mean that this generalizes and the advantage of increasing the number of predicted landmarks has been demonstrated for RGB images [62].
@@ Appendix F Societal Impact Statement
Figure 18: Example images from FAKE and T-FAKE with the augmentations applied during training.